OpenStudy (anonymous):

Tell me if I did this right Solve using the multiplication principle -5/6x=-7/8 answer: x=-21/20

OpenStudy (anonymous):

Yes your right. (cali love). http://www.youtube.com/watch?v=Py_0PtwJxzA

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

DEFINITELY SURE!

OpenStudy (phi):

Is this the problem? \[\frac{-5}{6x}=\frac{-7}{8}\] If so, cross multiply \[-5\cdot8= -7\cdot6x\] this simplifies to -40= -42x divide both sides by -42 to solve for x. of the problem is \[\frac{-5}{6}x= \frac{-7}{8}\] you get x= 42/40 or 21/20 (this almost matches your answer, but not quite)